OpenStudy (anonymous):

Add (−5t^2 − t + 3.5) + (5t^2 + 3t − 1.2)

OpenStudy (anonymous):

what do you think it is? it's very similar to the other problems you've posted

OpenStudy (anonymous):

i was thinking 3t + 2?

OpenStudy (anonymous):

not quite. the -5t^2 and 5t^2 cancel, as you have shown. you're left with - t + 3t and 3.5 - 1.2 -t is just (-1)t another way to think about it is (3-1) * t 3.5 - 1.2 is simple subtraction

OpenStudy (anonymous):

2t + 2.3?

OpenStudy (anonymous):

yep yep

OpenStudy (anonymous):

thank you!! you're so helpful

OpenStudy (anonymous):

do you know how to do this (−4a^2)(5a^5) sorry for asking you so many questions

OpenStudy (anonymous):

lol that's fine, and yeah: you can multiply exponentials together if the base is the same, for instance: 4^2 * 4^3 would be 4^5 keep the base the same, and add the exponents. (−4a^2)(5a^5) would be like: [-4 * 5] * [a^2 * a^5] you can pull out the numbers in front, and just look at the base and exponent. think you know it?

OpenStudy (anonymous):

i got -20^7 but when i put that it said its wrong because i had to simplify it but i'm not so good at simplifying

OpenStudy (anonymous):

-20a^7 you forgot to keep the a! unless that is a typo there is no way to further simplify it... unless they want like: -20a * -20a * -20a * -20a * -20a * -20a * -20a [-20a multiplied by itself 7 times, which is what an exponent is]

OpenStudy (anonymous):

ok i seee let me check because maybe i did forget the a

OpenStudy (anonymous):

yea you were right i forgot it

OpenStudy (anonymous):

haha it's all good

OpenStudy (anonymous):

its probably gonna be like 4 more questions i ask u if i can't get them lol if thats ok with you but what about this one 7b4(3b2 − 3b) when i did it i got 21b^6 - 21b idk

OpenStudy (anonymous):

first part is right, second part, not quite. 7b^4 * -3b is just -21 * b^(4+1) b is just shorthand for what it really means: b^1

OpenStudy (anonymous):

oooh ok didn't get that part so its 21b^5

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

well -21b^5, don't forget the negative!

OpenStudy (anonymous):

what about this one i have no idea what to do with it (x + 2)(x2 − 2x + 4) lol i feel like i'm asking too many questions

OpenStudy (anonymous):

well, to solve something like that, you need to distribute all the terms. it would be like: x * x^2 - x*2x + 4*x + 2*x^2 - 2*2x + 4*2 see what i did there?

OpenStudy (anonymous):

yes i see what you did but don't know how to solve it because its confusing

OpenStudy (anonymous):

well technically you cant solve this, as it doesn't equal anything. you are just un-factoring it. all you have to do is collect like terms it's just a lot of algebra

OpenStudy (anonymous):

yea it is which is really confusing lol

OpenStudy (anonymous):

try it and see what answer you get. its a bit confusing, but just go term-by-term and eventually after doing a couple problems you can do it in your head

OpenStudy (anonymous):

that x2 is supposed to be x^2

OpenStudy (anonymous):

that's what i figured, i wrote it that way when i distributed it all

OpenStudy (anonymous):

oh ok lol so the like factors would be 2 and 4? and idk what else

OpenStudy (anonymous):

another way of thinking about this is like this: (x + a) (x^2 + x + b) is the same as x * (x^2 + x + b) + a * (x^2 + x + b)

OpenStudy (anonymous):

well, if the question is: (x + 2)(x2 − 2x + 4) = 0 you can set x + 2 = 0 and x^2 - 2x + 4 = 0 x + 2 = 0 would juts be x = -2 however, you can't factor the second part, so you need to use the quadratic formula.

OpenStudy (anonymous):

oh ok i see. i'm probably not going to get this question lol

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

a = 1, b = -2, and c = 4

OpenStudy (anonymous):

not really is that the formula?

OpenStudy (anonymous):

no the formula is

OpenStudy (anonymous):